X+Y=10
X+Z=20
Y+Z=24
WHAT IS X+Y+Z

Answer by macston(5194)

X+Y=10
Y=10-X Use this to substitute for Y.
.
X+Z=20
Z=20-X Use this to substitute for Z.
.
Y+Z=24 Substitute for Y and Z
(10-X)+(20-X)=24
30-2X=24
-2X=-6
X=3
.
X+Y=10 Put in value for X.
3+Y=10
Y=7
.
X+Z=20 Plug in value for X.
3+Z=20
Z=17
.
CHECK:
Y+Z=24
7+17=24
24=24
.
X+Y+Z=?
3+7+17=?
27=?
ANSWER: The sum X+Y+Z=27